# Conformal Temporal Logic Planning using Large Language Models

### v4

@@ V Numerical Experiments
In this section, we provide extensive comparative experiments against existing LLM planners that require the task description in NL. In Section V-A, we compare HERACLEs against baselines that do not allow robots to ask for help [50, 81, 40]. These experiments show that the performance gap between these methods and HERACLEs increases significantly as task complexity increases even when the help-mode is disabled from our planner. In Section V-B, we compare our planner against KnowNo [41], a baseline that also employs CP allowing robots to ask for help in cases of high uncertainty. Our comparisons show that HERACLEs achieves significantly lower help rates. In that section, we also empirically validate the theoretical mission success rate guarantees discussed in Section IV. In Section V-C, we illustrate the ask-for-help mode of our planner on mobile manipulation tasks; see also [86]. In Section V-D, we demonstrate how various definitions of the NL-based predicates may affect performance of HERACLEs. Finally, in Section V-E, we provide examples to compare LTL and LTL-NL formulas in terms of their user-friendliness. In all case studies, we pick GPT-3.5, Llama 2-13b, and Llama 3-8b as the LLMs.
Model Method Case Study
@@
GPT 3.5 Ours $96\%$ $93.3\%$ $93\%$
SayCan [50] $96\%$ $40\%$ $14.08\%$
CMAS [81] $56\%$ $20\%$ $0\%$
ProgPrompt [40] $48\%$ $16\%$ $2.5\%$
Llama 2-13b Ours $86\%$ $81\%$ $80\%$
SayCan [50] $86\%$ $18\%$ $10\%$
CMAS [81] $53\%$ $9\%$ $0\%$
ProgPrompt [40] $46.7\%$ $9\%$ $0\%$
Llama 3-8b Ours $87.5\%$ $90.9\%$ $90\%$
SayCan [50] $87.5\%$ $36\%$ $18\%$
CMAS [81] $50\%$ $27\%$ $18\%$
ProgPrompt [40] $43.8\%$ $18.1\%$ $0\%$
@@
TABLE I: Comparison of planning accuracy of HERACLEs (when the CP/help mode is deactivated) against non-conformalized baselines, using GPT 3.5, Llama 2-13b, and Llama 3-8b.
V-A Comparisons against Non-Conformalized Neural Planners
[…]
Figure 3: Case Study I: Execution of a robot plan for the LTL-NL mission: $\Diamond$ (’Move Coke 1 to Location B’) requiring the robot to eventually deliver Coke 1 to location B. The left snapshot illustrates the robot plan generated by HERACLEs. The right snapshot shows the environment when the task is completed.
Figure 4: Case Study II: Execution of a robot plan for the LTL-NL mission: $\phi=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge(\neg\pi_{1}{\mathcal{U}}\pi_{2})$, where $\pi_{1}$ refers to ‘Deliver Apple to A’ and $\pi_{2}$ refers to ‘Deliver Coke 2 to A’. Notice here $\phi$ requires that the apple should not be delivered to Location A prior to Coke 2. The left two snapshots illustrate the sub-tasks that the robot accomplished in the order that they were generated by the symbolic planner. The corresponding robot plans are also depicted. The right snapshot shows the environment when the task is completed.
Baselines: As the baselines for our experiments, we employ three state-of-the-art LLM-based planners: SayCan [50], CMAS [81], and ProgPrompt[40]. Unlike HERACLEs, these baselines require the overall mission to be fully described using NL. Thus, we manually convert LTL-NL tasks into NL ones, which are then used as inputs for the baselines. SayCan generates decisions, taking into consideration the likelihood that the corresponding action will be executed correctly. Since we consider perfect execution of robot actions, we assume that the affordance functions in [50] always return a value of $1$ for all decisions. SayCan converts the NL mission planning problem into a sequence of MCQA problems, as HERACLEs does too. Then SayCan selects the action with the highest softmax score: $s(t)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t))$ for every MCQA problem. We also compare our method against the CMAS planner [81]. CMAS is originally developed for teams of $N$ robots. We applied it in our setting using $N=1$. Note that CMAS and ProgPrompt do not employ the MCQA setup (unlike HERACLEs and SayCan). Instead, at every time step, the LLM generates new tokens corresponding to robot actions. ProgPrompt - unlike our method, Saycan, and CMAS - requires its prompt to be written in the format of an executable programming language such as Python. We compare the planners based on their runtimes as well as their accuracy over scenarios drawn from ${\mathcal{D}}$ defined as the percentage of scenarios where a planner generates a correct plan; see Tables I-II.
To make our comparisons fair, we have enforced the following requirements: (i) All methods are exposed to the same set of actions ${\mathcal{S}}$. (ii) All methods (except ProgPrompt) share the same prompt structure. The only difference is that part (C) in the prompts for HERACLEs includes the description of the sub-task generated by the symbolic planner while part (C) for SayCan, CMAS, and ProgPrompt includes the overall mission expressed in NL (as these baselines do not consider task decomposition). The main body of the prompt in ProgPrompt is described in Python-like format as in the original work [40]. (iii) All actions are executed perfectly. (iv) We completely deactivate the help mode from our planner since this capability does not exist in these baselines. We emphasize that we have applied (iv) only to make comparisons fair against the baselines as otherwise our planner can outperform them by picking a low enough value for $\alpha$ and asking for help from users. This choice also allows us to assess the ‘nominal’ performance of our planner when help from users is not available; evaluation of our method without removing CP is presented in Sections V-B. The requirement (iv) means that we remove altogether the CP component from our planner. This implies that HERACLEs always picks the decision $s(t)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t))$ (as SayCan does too). Also, to accommodate (iv), we consider environments with known obstacles that do not prevent the robot from accomplishing any sub-task included in $\phi$. This implies that HERACLEs will never trigger an assistance request from the symbolic planner. We note again that have applied (iv) only to ensure fairness to other baselines that do not have the capability of asking for help from users or symbolic planners; demonstrations in environments that prevent the accomplishment of certain sub-tasks are provided in Section V-C. (v) We require all methods to complete the plan within $H=T\times K$ steps, where $K$ is the number of predicates in $\phi$ and $T=7$ is the hyper-parameter used in Section III-C.
In what follows, we sample $111$ scenarios from ${\mathcal{D}}$ and group them based on their difficulty category. We first discuss the planning accuracy of all methods when they are coupled with GPT-3.5 and Llama 2-13b. These results are summarized in Table I which also includes the performance of the methods when equipped with Llama 3-8b.
@@ V-A Comparisons against Non-Conformalized Neural Planners
Case Study I (Easy): We consider $25$ LTL-NL formulas of the form $\phi=\Diamond\pi_{1}$ where $\pi_{1}$ is defined as ‘Move object $e$ to location ${\mathbf{x}}$” for various objects $o_{e}\in{\mathcal{O}}$ and locations ${\mathbf{x}}$. We manually translate such formulas into NL as ‘Eventually move object $e$ to location ${\mathbf{x}}$’ to use as the task specification for CMAS and SayCan. Using GPT 3.5, the accuracy of HERACLEs was $96\%$ while the accuracy of SayCan, CMAS, and ProgPrompt were $96\%$, $56\%$, and $48\%$, respectively. Using Llama 2-13b, the accuracy of our method and SayCan dropped to 86%, while for CMAS and ProgPrompt, it dropped to 53% and 46.7%, respectively. This accuracy drop is expected as Llama 2-13b is a smaller model compared to GPT 3.5. Notice that the performance of our method and SayCan is the same since the mission $\phi$ cannot be further decomposed and both methods share the same prompt as well as the same action selection mechanism when the help mode is removed from our method. A robot plan generated by HERACLEs using GPT 3.5 is shown in Fig. 3.

### v5

@@ 5. Experiments
In this section, we demonstrate the performance of the proposed method in various numerical and hardware experiments. First, in Section 5.1, we empirically validate the mission success rate guarantees of HERACLEs (see Theorem 4.2). We also compare our method against KnowNo (Ren et al., 2023). We select this baseline as it can also achieve desired mission success rates by letting robots ask for help, using CP, in cases of high uncertainty. Second, in Section 5.2 we provide hardware experiments on real robot platforms to test how the mission completion rate guarantees of HERACLEs may be affected by unmodeled sources of uncertainty (e.g., imperfect robot skills). Demonstrations of HERACLEs are provided in (Demonstrations, 2024). Third, in Section 5.3, we provide examples to compare LTL and LTL-NL formulas in terms of their user-friendliness. Finally, in Section 5.4, we demonstrate how various definitions of the NL-based predicates can affect performance of HERACLEs. Comparisons against baselines that do not allow robots to ask for help are presented in Appendix A. In all case studies, we pick the following LLMs: GPT-3.5, Llama 2-13b, Llama 3-8b, and Qwen 3-32b.
5.1. Comparative Numerical Experiments
In this section, we empirically validate the mission success rate guarantees of Theorem 4.2 on mobile manipulation tasks. We also provide comparative experiments against KnowNo (Ren et al., 2023), in terms of help rates required to achieve desired task success rates as well as length of plans and runtimes. Overall, HERACLEs achieves significantly lower help rates due to its symbolic component.
Environment & Robot Setup: We consider mobile manipulation tasks over an environment populated with $6$ objects: two cans of Coke, one water bottle, one pen, one tin can, and one apple. The water bottle is inside the fridge and the pen is inside a drawer. Thus, grabbing e.g., the pen requires the robot to first open the drawer if it is closed. The status of these containers (open or closed) is not known a-priori and, therefore, not included in the initial environment description in $\ell(1)$. Instead, it can be provided online through sensor feedback as described in Section 3.3. The action space $\mathcal{A}$} includes $5$ actions as defined in Fig. 4. The action ‘remain idle’ in $\mathcal{A}$} is useful when a sub-task can be accomplished in less than $T$ time steps (see Section 3.3). Given a prompt $\ell(t)$, the number of choices $s$ that the LLM can pick from is $|{\mathcal{S}}|=17$. Recall that this set is constructed using $\mathcal{A}$} and all objects/locations in the environment where the actions in $\mathcal{A}$} can be applied. We select $T=7$ for the LLM-based action planner.
LTL-NL Tasks: We construct a distribution $\mathcal{D}$ to sample LTL-NL formulas following the approach in (Ren et al., 2023). For all generated formulas, there exists only one sequence of sub-tasks that can lead to mission completion as required by Th. 4.2 and Appendix B; this is relaxed in Section 5.2. Specifically, $\mathcal{D}$ is designed to generate scenarios $\xi$ of ‘easy,’ ‘medium,’ and ‘hard’ difficulty, with probabilities $0.2$, $0.1$, and $0.7$, respectively. Each difficulty category comprises hundreds of LTL-NL formulas, with the difficulty level determined by the complexity of the formulas in terms of the number of NL-based predicates and temporal/logical operators. We note that this may not be an accurate classification of the mission scenarios as task complexity may depend on additional parameters (e.g., number of actions required for completion of each predicate/sub-task or whether the sub-tasks requiring opening containers to retrieve items) while also what makes a task complex may really depend on the employed LLM. When a difficulty category is chosen, a random LTL-NL formula is selected from the corresponding predefined set. We emphasize that $\mathcal{D}$} is not known to our framework. More complex distributions $\mathcal{D}$} can also be constructed as CP, and, consequently, HERACLEs, is distribution-free.
We sample $111$ scenarios from $\mathcal{D}$} and group them based on their difficulty category. Easy tasks include $25$ LTL-NL formulas of the form $\phi=\Diamond\pi_{1}$ where $\pi_{1}$ is defined as ‘Move object $e$ to location $\mathbf{x}$}’ for various objects $e$ and locations $\mathbf{x}$}; Medium tasks include $15$ LTL-NL formulas defined as $\phi=\Diamond\pi_{1}\wedge(\neg\pi_{1}{\mathcal{U}}\pi_{2})$, such a task requires eventually completing the sub-tasks $\pi_{1}$ and $\pi_{2}$, but $\pi_{2}$ must be completed strictly before $\pi_{1}$. Hard tasks include $71$ LTL-NL formulas defined over four to six predicates defined as before. An example of such formula is $\phi=(\neg(\pi_{1}\vee\pi_{2}\vee\pi_{3}\vee\pi_{4})){\mathcal{U}}(\pi_{1}\wedge\neg\pi_{2}\wedge\neg\pi_{3}\wedge\neg\pi_{4}\wedge\rightarrow(\neg(\pi_{1}\vee\pi_{2}\vee\pi_{3}\vee\pi_{4}){\mathcal{U}}(\pi_{2}\wedge\neg\pi_{1}\wedge\neg\pi_{3}\wedge\neg\pi_{4}\wedge\rightarrow(\neg(\pi_{1}\vee\pi_{2}\vee\pi_{3}\vee\pi_{4}){\mathcal{U}}(\pi_{3}\wedge\neg\pi_{1}\wedge\neg\pi_{2}\wedge\neg\pi_{4}\wedge\rightarrow(\neg(\pi_{1}\vee\pi_{2}\vee\pi_{3}\vee\pi_{4}){\mathcal{U}}(\pi_{4}\wedge\neg\pi_{1}\wedge\neg\pi_{2}\wedge\neg\pi_{3})))))))$, which enforces that $\pi_{1}$, $\pi_{2}$, $\pi_{3}$, and $\pi_{4}$ must be done only once, and in this specific order exclusively.
Setting Up the Baseline: To enable meaningful and fair comparisons against the KnowNo baseline, we have considered the following settings. First, since KnowNo uses NL for mission specification, we convert the LTL-NL formulas into NL instructions by translating each temporal/logical operator and atomic proposition into their corresponding NL description. For instance, a medium-level task $\phi=\Diamond\pi_{1}\wedge(\neg\pi_{1}{\mathcal{U}}\pi_{2})$ can be translated as ‘Eventually finish {NL description of $\pi_{1}$} and {NL description of $\pi_{2}$}. Do not do {NL description of $\pi_{1}$} until {NL description of $\pi_{2}$} is completed’. The resulting NL instructions serve as an input to KnowNo. Second we set up the MCQA-based prompt of KnowNo using the set of decisions/options $\mathcal{S}$} that HERACLEs uses; see also Remark 5.2. Third, we consider fully known environments (e.g., known obstacles) that do not prevent the robot from accomplishing any sub-task included in $\phi$. As a result, both HERACLEs and KnowNo are applied offline to compute plans. We emphasize that we have considered this setting only to ensure fairness to the baseline. The reason is that our method allows the robot to seek assistance from the symbolic planner in case the environment prevents completion of a sub-task; a feature that does not exist in KnowNo. Demonstrations in environments that prevent the accomplishment of certain sub-tasks are presented in (Demonstrations, 2024) and discussed in Section 5.2. Fourth, we require KnowNo to complete the plan within $H=T\times K$ steps, where $K$ is the number of predicates in $\phi$ and $T=7$ is the hyper-parameter used in Section 5.2. We made this choice since HERACLEs also completes a sub-plan for each sub-task in $\phi$ within $T$ steps.
Model Method $1-\alpha=0.90$ $1-\alpha=0.95$
Success Help Success Help
Llama 2-13b Ours 93.3% 9% 97.2% 12%
KnowNo 93.6% 67.2% 96.7% 78%
Llama 3-8b Ours 91.0% 20% 98.1% 24%
KnowNo 90.3% 55.8% 95.1% 64.4%
Qwen 3-32b Ours 92.9% 8.4% 97.7% 10.1%
KnowNo 90.5% 53.2% 95.4% 61.6%
Table 1. Comparison of mission success and user help rates for HERACLEs and KnowNo across Llama2-13b, Llama3-8b, and Qwen 3-32b under two confidence levels.
Empirical Mission Success Rates & Help Rates for HERACLEs: We empirically validate the theoretical guarantees for the mission success rate of HERACLEs (Theorem 4.2) using Llama 2-13b, Llama 3-8b, and Qwen 3-32b models. For each validation scenario, we collect $20$ calibration sequences from $\mathcal{D}$}. A plan is then computed using Alg. 1, while allowing the robot to request help from users when needed. The empirical mission success rate is obtained as follows: among the $111$ generated plans, we compute the ratio of those that are ground truth with respect to their corresponding LTL-NL task $\phi$. This process is repeated $100$ times, and the average ratio is reported as the empirical mission success rate. Table 1 summarizes the results for two confidence levels ($1-\alpha=0.90$ and $1-\alpha=0.95$). At $1-\alpha=0.90$, HERACLEs achieves success rates of $93.3\%$, $91.0\%$, and $92.9\%$ with Llama 2-13b, Llama 3-8b, and Qwen 3-32b, respectively. These validate the theoretical guarantees. The corresponding help rates are $9\%$, $20\%$, and $8.4\%$. At $1-\alpha=0.95$, the success rates further increase to $97.2\%$, $98.1\%$, and $97.7\%$, respectively, with help rates of $12\%$, $24\%$, and $10.1\%$. Observe that, as expected, the help rates increase as $1-\alpha$ increases. A robot plan synthesized for one of the considered medium tasks using HERACLEs and GPT 3.5 is shown in Fig. 6
Figure 6. The execution of a robot plan using ROS/Gazebo (Marian et al., 2020) for the LTL-NL mission: $\phi=\Diamond\pi_{1}\wedge(\neg\pi_{1}{\mathcal{U}}\pi_{2})$, where $\pi_{1}$ refers to ‘Deliver Apple to A’ and $\pi_{2}$ refers to ‘Deliver Coke 2 to A’. Here $\phi$ requires that the apple should not be delivered to A prior to Coke 2. The left two snapshots illustrate the NL-based sub-tasks that the robot accomplished in the order that they were generated by the symbolic planner. The corresponding robot plans are also depicted. The right snapshot shows the environment when the task is completed.
Comparisons against KnowNo: We repeated the above process for KnowNo equipped with all three models, using the same validation and calibration data as HERACLEs. Specifically, with Llama 3-8b, KnowNo achieves $90.25\%$ and $95.12\%$ mission success rates with help rates of $55.8\%$ and $64.4\%$ at $1-\alpha=0.9$ and $1-\alpha=0.95$, respectively. With Llama 2-13b, KnowNo achieves $93.6\%$ and $96.7\%$ success rates, with help rates of $67.2\%$ and $78\%$. With Qwen 3-32b, KnowNo achieves $90.5\%$ and $95.4\%$ success rates, but requires help rates of $53.2\%$ and $61.6\%$, respectively. Overall, both methods satisfy the theoretical guarantees on mission success rates. However, KnowNo consistently demands much higher help rates across all models. This highlights the benefit of the symbolic component in HERACLEs, which decomposes the planning problem into smaller sub-problems that the LLM can solve more effectively. In contrast, KnowNo directly relies on the LLM to plan for the entire long-horizon task. In terms of runtime, both methods are comparable: generating a complete plan takes approximately $70$ seconds with Llama 2-13b, $75$ seconds with Llama 3-8b, and $288$ seconds with Qwen 3-32b. [^3] We note that these runtimes are implementation-specific. In our implementation, to design an action $s(t)$, using either method, we query the LLM sequentially, $|{\mathcal{S}}|$ times, to obtain the confidence score per option. These scores are used to construct $\mathcal{C}(\ell(t))$. These $|{\mathcal{S}}|$ queries at time $t$ could be conducted in parallel, since they are independent of one another, to accelerate the construction of $\mathcal{C}(\ell(t))$; however, this was not possible in our implementation due to limited memory. Each query required about $0.4$ seconds on average with the Llama models, and about $1.6$ seconds with Qwen 3-32b. The LTL planner in HERACLEs further requires only $2.75\times 10^{-5}$ seconds on average to generate sub-tasks. Finally, the average length of successful plans across all difficulty levels is $15$ steps for both methods, excluding idle actions.
Remark 5.1 (LLM model & Help Rates).
The quality of the LLM model can critically affect the help rates, with more effective models resulting in lower rates. For instance, the help rate of our method with GPT 3.5 is $2.56\%$, $2.63\%$, and $2.70\%$ for $1-\alpha=0.95$, $0.96$, and $0.97$, respectively.
Remark 5.2 (Set of Options for KnowNo).
KnowNo, as originally designed in (Ren et al., 2023), queries an LLM at each time step to generate a finite set of robot decisions, which are presented as choices in its MCQA-based prompt. These choices include an additional option ‘other option, not listed’ to account for cases where all LLM-generated options are nonsensical. If the LLM fails to provide valid options, then that would yield an ‘incomplete’ plan ending with ‘other option, not listed’, which would make direct comparisons non-trivial. Thus, we expose KnowNo to the set $S$ of decisions that HERACLES uses, without asking the LLM to generate a set of options.
5.2. Hardware Experiments
In this section, we conduct hardware experiments to evaluate robustness of the proposed method against un-modeled sources of uncertainty such as imperfect execution of robot skills.
Environment & Robot Setup: We consider a robot tasked with road-maintenance missions, in a post-disaster city, requiring the robot to block certain streets by placing traffic cones and remove damaged fire hydrants in a temporal/logical order. The environment is populated with $4$ traffic cones and $2$ fire hydrants. Unlike the numerical experiments in Section 5.1, here the geometric structure of the environment is unknown and it may prevent access to certain objects of interest. The action space $\mathcal{A}$} is the same as in Section 5.1 except the action ‘$(4,x)$’ is removed as there are no containers in the considered environment. This results in a set $\mathcal{S}$} of $13$ decisions.
Empirical Mission Success Rates & Help Rates: We construct a distribution $\mathcal{D}$ designed to generate tasks of ‘easy’, ‘medium’, and ‘hard’ difficulty levels—similar to those in Section 5.1—but which may admit multiple feasible sequences of sub-tasks; see, for example, the LTL formula in Fig. 1. We generate $30$ validation scenarios from $\mathcal{D}$} and for each validation scenario, we collect $15$ calibration scenarios. Using HERACLEs, coupled with Llama 3-8B, we compute robot plans for $1-\alpha=0.9$. The mission success rate is $96.7\%$ with the help rate of $6.3\%$.
Execution of Plans: We select $12$ validation scenarios where HERACLEs successfully designed a correct plan and tested them on a ground robot platform. For these experiments, we use a TurtleBot3 Waffle Pi equipped with a 4-DOF OpenManipulator-X (OpenRobotics, 2022) operating within the Mini-City platform developed at WashU. The robot employs existing SLAM algorithms for state estimation and mapping (Grisetti et al., 2007). In $11$ out of $12$ scenarios, the robot successfully executed the plans without any skill failures. The single failed scenario occurred due to a traffic cone slipping off the robot manipulator. Snapshots of the robot navigating the Mini-City to satisfy the LTL-NL formula used in Example 2.3, while also asking for help, are provided in Fig. 1. Additional demonstrations of our hardware experiments are provided in (Demonstrations, 2024). Future work will focus on designing robust planners to address uncertainties encountered during deployment.
5.3. Comparisons of LTL-NL and LTL Formulas
In this section, we provide examples to demonstrate that LTL-NL formulas are more user-friendly than their corresponding LTL formulas defined over the system states ${\mathbf{p}}(t)$ capturing the same mission requirements. As a performance metric, we employ the ‘length’ of the specification, i.e., the total number of atomic predicates and temporal and logical operators. In what follows, we select an LTL-NL formula for each case study considered in Appendix A and we manually convert it into the respective LTL formula. First, consider the LTL-NL formula $\phi=\Diamond\pi_{1}$ from Case Study I, where the NL-based predicate $\pi_{1}$ is true if the robot delivers a bottle of water to location $A$. This formula only has $1$ temporal/logical operator and $1$ predicate. Using system-based predicates, the same task will be written where $\pi_{1}^{\prime}$ is true if the robot position is close enough to the bottle of water, $\pi_{2}^{\prime}$ is true if the robot grabs the bottle successfully, $\pi_{3}^{\prime}$ is true if the robot position is close enough to location $A$, and $\pi_{4}^{\prime}$ is true if the robot puts down the bottle of water. Second, consider the LTL-NL formula $\phi=\Diamond\pi_{1}\wedge\Diamond\pi_{2}$ from Case Study II where both $\pi_{1}$ and $\pi_{2}$ model delivery tasks as before. This formula has $3$ logical/temporal operators and $2$ predicates. Expressing the same task using system-based predicates as before would result in $15$ logical/temporal operators and $8$ predicates, i.e., $\phi=\Diamond(\pi_{1}^{\prime}\wedge(\Diamond\pi_{2}^{\prime}\wedge(\Diamond\pi_{3}^{\prime}\wedge(\Diamond\pi_{4}^{\prime}))))\wedge\Diamond(\pi_{5}^{\prime}\wedge(\Diamond\pi_{6}^{\prime}\wedge(\Diamond\pi_{7}^{\prime}\wedge(\Diamond\pi_{8}^{\prime}))))$. Third, we select the following formula from Case Study III: $\phi=\Diamond\pi_{1}\wedge\Diamond\pi_{2}\wedge\Diamond\pi_{3}\wedge\Diamond\pi_{4}\wedge(\neg\pi_{4}{\mathcal{U}}\pi_{1})$. This LTL-NL formula has $10$ temporal/logical operators and $4$ predicates, whereas the corresponding LTL formula over system-based predicates requires $46$ temporal/logical operators and $16$ predicates. Observe that the difference in the length of LTL and LTL-NL formulas becomes more pronounced as the complexity of the mission requirements increases.
5.4. Effect of Predicate Definitions on Planning Performance
Specifically, we consider a simple mission requiring to deliver objects A, B, C, and D to desired destinations in any order. We can express this task using an LTL-NL formula with four APs: $\phi_{1}=\Diamond\pi_{A}\wedge\Diamond\pi_{B}\wedge\Diamond\pi_{C}\wedge\Diamond\pi_{D}$. In $\phi_{1}$, $\pi_{i}$ is true when the object $i\in\{A,B,C,D\}$ is delivered to the intended destination. An equivalent LTL-NL formula is $\phi_{2}=\Diamond\pi_{A,B}\wedge\Diamond\pi_{C}\wedge\Diamond\pi_{D}$ where the only difference is that now we have defined a predicate $\pi_{A,B}$ that is true if both $A$ and $B$ are delivered to their destinations in any order. In a similar way, we can define the equivalent LTL-NL formulas $\phi_{3}=\Diamond\pi_{A,B,C}\wedge\Diamond\pi_{D}$ and $\phi_{4}=\Diamond\pi_{A,B,C,D}$. Observe that formula $\phi_{4}$ is the most ‘user-friendly’ one as it has the fewest predicates and temporal operators. However, this results in increasing the number of sub-tasks that each predicate models. This in turn requires the LLM to reason about more complex task requirements that would otherwise have been handled by the symbolic planner. For instance, in $\phi_{4}$, the symbolic planner generates the subtask $\pi_{A,B,C,D}$ and the LLM needs to design a feasible plan for it. Increasing the complexity of the sub-tasks modeled by the predicates may in turn result in higher help rates. Specifically, we run HERACLEs using Llama 2-13b model to compute plans for each LTL-NL formula. When $1-\alpha=0.9$, the average help rates over $10$ trials for $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, and $\phi_{4}$ were $10\%$, $35\%$, $50\%$, and $85\%$ respectively; each trial differs in the calibration dataset. Notice that for $\phi_{4}$, HERACLEs is essentially equivalent to (Ren et al., 2023) since the symbolic planner cannot perform task decomposition by construction of $\phi_{4}$.
6. Conclusions, Limitations, and Future Work
In this paper, we proposed HERACLEs, a new neuro-symbolic planner for LTL-NL tasks. We showed, both theoretically and empirically, that it can achieve desired mission success rates due to a conformal interface between the symbolic planner and the LLM. We also provided comparative experiments showing that it outperforms LLM-based planners in terms of planning accuracy and help rates as well as examples to demonstrate its user-friendliness compared to symbolic planners. Future work will address imperfect skill execution and out-of-distribution mission scenarios.
Appendix A Comparisons against Non-Conformalized Neural Planners
In this Appendix, we compare HERACLEs against baselines (Ahn et al., 2022; Chen et al., 2023b; Singh et al., 2023) which, however, do not utilize CP and do not allow robots to ask for help. Our comparative experiments show that the ability of these approaches to design correct plans tends to decrease as the number of temporal and logical requirements in the mission increases. On the contrary, the planning performance of HERACLEs consistently outperforms these baselines (even if the help mode is deactivated). We consider the same environment and robot setup as in Section 5.1.
Setting up the Baselines: As the baselines for our experiments, we employ three state-of-the-art LLM-based planners: SayCan (Ahn et al., 2022), CMAS (Chen et al., 2023b), and ProgPrompt(Singh et al., 2023). Unlike HERACLEs, these baselines require the overall mission to be fully described using NL. Thus, we manually convert LTL-NL tasks into NL ones, which are then used as inputs for the baselines; see Section 5.1. SayCan generates decisions, taking into consideration the likelihood that the corresponding action will be executed correctly. Since we consider perfect execution of robot actions, we assume that the affordance functions in (Ahn et al., 2022) always return a value of $1$ for all decisions. We note that SayCan converts the NL mission planning problem into a sequence of MCQA problems, as HERACLEs does too. Then SayCan selects the action with the highest softmax score: $s(t)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t))$ for every MCQA problem. We also compare our method against the CMAS planner (Chen et al., 2023b). CMAS is originally developed for teams of $N$ robots. We applied it in our setting using $N=1$. Note that CMAS and ProgPrompt do not employ the MCQA setup (unlike HERACLEs and SayCan). Instead, at every time step, the LLM generates new tokens corresponding to robot actions. ProgPrompt - unlike our method, Saycan, and CMAS - requires its prompt to be written in the format of an executable programming language such as Python.
To make our comparisons fair, we have enforced the following requirements: (i) All methods are exposed to the same set of actions $\mathcal{S}$}. (ii) All methods (except ProgPrompt) share the same prompt structure. The only difference is that part (C) in the prompts for HERACLEs includes the description of the sub-task generated by the symbolic planner while part (C) for SayCan, CMAS, and ProgPrompt includes the overall mission expressed in NL (as these baselines do not consider task decomposition). The main body of the prompt in ProgPrompt is described in Python-like format as in the original work (Singh et al., 2023). (iii) All actions are executed perfectly. (iv) We completely deactivate the help mode from our planner since this capability does not exist in these baselines. We emphasize that we have applied (iv) only to make comparisons fair against the baselines as otherwise our planner can outperform them by picking a low enough value for $\alpha$ and asking for help from users. This choice also allows us to assess the ‘nominal’ performance of our planner when help from users is not available. The requirement (iv) means that we remove altogether the CP component from our planner. This implies that HERACLEs always picks the decision $s(t)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t))$ (as SayCan does too). Also, to accommodate (iv), we consider environments with known obstacles that do not prevent the robot from accomplishing any sub-task included in $\phi$; see Section 5.1. This implies that HERACLEs will not trigger an assistance request from the symbolic planner. Again, we note that we have applied (iv) only to ensure fairness to other baselines that do not have the capability of asking for help from users or symbolic planners. (v) We require all methods to complete the plan within $H=T\times K$ steps, where $K$ is the number of predicates in $\phi$ and $T=7$ is the hyper-parameter used in Section 3.3. In what follows, we report the performance of our method and the baselines over the LTL-NL formulas of various levels of difficulty. As performance metrics, we use runtimes as well as planning accuracy defined as the percentage of scenarios where a planner generates a correct plan. The results are also summarized in Table 2.
Model Method Case Study Avg. Runtime (s)
@@ Appendix A Comparisons against Non-Conformalized Neural Planners
GPT 3.5 Ours 96% 93.3% 93% 14
SayCan $96\%$ $40\%$ $14.08\%$ 15
CMAS $56\%$ $20\%$ $0\%$ 3
ProgPrompt 48% 16% 2.5% 4
Llama 2-13b Ours 86% 81% 80% 71
SayCan $86\%$ $18\%$ $10\%$ 75
CMAS $53\%$ $9\%$ $0\%$ 15
ProgPrompt 46.7% 9% 0% 18
Llama 3-8b Ours 87.5% 90.9% 90% 76
SayCan $87.5\%$ $36\%$ $18\%$ 78
CMAS $50\%$ $27\%$ $18\%$ 17
ProgPrompt 43.8% 18.1% 0% 25
Qwen 3-32b Ours 96% 93.3% 95% 300
SayCan 96% 63% 25% 350
CMAS 80% 48% 30% 72
ProgPrompt 96% 93% 42% 90
@@
Table 2. Comparison of planning accuracy and runtime for HERACLES (CP/help mode deactivated) against non-conformalized baselines, using GPT 3.5, Llama 2-13b, Llama 3-8b, and Qwen 3-32b.
@@
Case Study I (Easy): We consider $25$ LTL-NL formulas of the form $\phi=\Diamond\pi_{1}$ where $\pi_{1}$ is defined as ‘Move object $e$ to location $\mathbf{x}$}’ for various objects $e$ and locations $\mathbf{x}$}. Using GPT 3.5, the accuracy of HERACLEs was $96\%$ while the accuracy of SayCan, CMAS, and ProgPrompt were $96\%$, $56\%$, and $48\%$, respectively. Using Llama 2-13b, the accuracy of our method and SayCan dropped to 86%, while for CMAS and ProgPrompt, it dropped to 53% and 46.7%, respectively. Using Qwen 3-32b, HERACLEs achieved 96% of accuracy, while SayCan, CMAS, and ProgPrompt achieved 96%, 80%, and 96%, respectively. Notice that the performance of our method and SayCan is the same since the mission $\phi$ cannot be further decomposed, and both methods share the same prompt as well as the same action selection mechanism when the help mode is removed from our method.